OpenStudy (anonymous):

y = log x If y = 10, then what is x?

OpenStudy (campbell_st):

x is 10^10

OpenStudy (anonymous):

My teacher says the answer is 1, but I dont know how she got that.

OpenStudy (anonymous):

\[\log(x)=10\] or \[\log(10)=x\]?

OpenStudy (anonymous):

what is the base of the logarithm here?

OpenStudy (anonymous):

because if \[\log(x)=10\] this means \[x=10^{10}\] (assuming base is 10)

OpenStudy (campbell_st):

this question uses exponentials.. \[10^y = 0^{\log x}\] then by log laws \[10^y = x\] if y = 10 then \[10^10 = x\]